OpenStudy (anonymous):

f(x) = 2x2 – 2x + 3 FIND THE ZEROS BY USING THE QUADRATIC FORMULA

OpenStudy (anonymous):

holy are you just posting every single one of your homework questions on this board

OpenStudy (anonymous):

\[\large f(x)=ax^2-bx+c\] \[\large f(x) = 2x^2 – 2x + 3\] Using the quadratic formula: \[\large x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] \[\large x=\frac{-(-2)\pm \sqrt{(-2)^2-4(2)(3)}}{2(2)}\] \[\large x=\frac{2\pm \sqrt{4-8(3)}}{4}\]

OpenStudy (anonymous):

You can tell me what the zeroes are. I gave you a headstart, now you can finish it off.

OpenStudy (anonymous):

COULD YOU HELP ME W/ANOTHER PROBLEM @Azteck

OpenStudy (anonymous):

If it's something to do with quadratics, it's best that you try it yourself first.

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

@Azteck

OpenStudy (anonymous):